OpenStudy (anonymous):

what is the relationship between a and c? a=b^(2/3) and b=c^(3/4)

OpenStudy (anonymous):

b^(2/3) = [c^(3/4)]^(2/3) = c^(1/2)

OpenStudy (anonymous):

could you explain a little further please?

OpenStudy (anonymous):

i just raised the second equation to the power of 2/3

OpenStudy (anonymous):

which gives b^(2/3) = c^(1/2) since a = b^(2/3), a = c^(1/2)

OpenStudy (anonymous):

oh, okay! Thanks!!

OpenStudy (jdoe0001):

keep in mind that \(\Large \bf a^{\frac{{\color{blue} n}}{{\color{red} m}}} = \sqrt[{\color{red} m}]{a^{\color{blue} n}}\)